MadisonMarvel:

What is the following product? StartRoot 30 EndRoot times StartRoot 10 EndRoot

michaellopez:

Simplify the radical by breaking the radicand up into a product of known factors, assuming positive real numbers.

michaellopez:

this is close to the answer 17.32050807

michaellopez:

now change it to its exact form

michaellopez:

@MadisonMarvel

xxcoffee:

Dont ask for the "best response" you can get in trouble for it

xxcoffee:

You may help them work to the anwser but do not ever ask for the ""best response" ik people that get ban or suspended for that

xxcoffee:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @michaellopez yah but i answered it and she should reward me for it \(\color{#0cbb34}{\text{End of Quote}}\) and others do the same thing but they dont ask for the "best response"

xxcoffee:

@dude

michaellopez:

why are you tagging him

xxcoffee:

so he can close it

michaellopez:

ok

xxcoffee:

@Ultrilliam

AZ:

\(\sqrt{a} \times \sqrt{b} = \sqrt{a \times b}\) (and remember, both a and b have to be positive numbers for you to use that rule) so what is \(\sqrt{30} \times \sqrt{10} = \)

AZ:

after that, we can simplify it :)